OpenStudy (anonymous):

At a local fitness center members pay a $12 dollar membership fee and $3 dollars for each aerobic class. Non members pay $5 dollars for each aerobic class. For what number of aerobic classes will the cost for members and non members be the same?

OpenStudy (anonymous):

6 classes

OpenStudy (anonymous):

Seriously? Thanks!

OpenStudy (anonymous):

12+3(x)=5(x) 12=5x-3x 12=2x 6=x

OpenStudy (anonymous):

Thank you! Awesome!

OpenStudy (anonymous):

no problem @thisisarri